OpenStudy (anonymous):

Determine which ordered pair is NOT a solution of y=-4x-2 a. (–5, 21) b. (3, –14) c. (–1, 2) d. (–9, 34) help, please.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Ash, the easiest to determine the answer of your question would be to plug in the values of x and y in the equation to see whether the left hand side equal right hand side. Doing that, you can check for yourself that a is the only one that contradicts the given equation.

OpenStudy (anonymous):

because when u put the value of x in the equation u don't get the value of y y=-4(-5)-2=20-2=18 and 18 doesn't equal 21